(9) M. Csoergoe, Gaussian processes, strong approximations: An interplay, in: Aspects statistiques et aspects physiques des processus gaussiens, Saint- Flour 1980, Colloq. Int. C.N.R.S. 307, 131-229 (1981).

(11) L. Pace, In: 14-th European Meeting of Statisticians, Abstracts, p. 244, Wrocław (1981).

(21) J. Sunklodas, Approximation of distributions of sums of weakly dependent random variables by the normal distribution. (Russian) R. V. Gamkrelidze (ed.) et al., Probability theory - 6. Limit theorems in probability theory. Moskva: Vsesoyuznyj Institut Nauchnoj i Tekhnicheskoj Informatsii, Itogi Nauki Tekh., Ser. Sovrem. Probl. Mat., Fundam. Napravleniya. 81, 140-199 (1991).

(156) Simeon Reich - David Shoikhet, The existence of resolvents of holomorphic generators in Banach spaces. (English) [A] Athanassios G. Kartsatos (ed.), Theory and applications of nonlinear operators of accretive and monotone type. New York, NY: Marcel Dekker. Lect. Notes Pure Appl. Math. 178, 251-258; p. 257 (1996).

(157) A C Payatakes, G N Constantinides, M S Valavanides; Hierarchical theoretical models: An informal introduction In: Mathematical methods in scattering theory and biomedical technology. 1997.06.30-1997.07.01. (1998) pp. 158-169. Zbl: G. Dassios (ed.) et al. Proceedings of a workshop dedicated to Professor Gary Roach, Harlow: Longman. Pitman Res. Notes Math. Ser. 390, 158-169; p. 158 (1998)

(170) A. Goldvard - S. Reich - D. Shoikhet; ASYMPTOTIC REPRESENTATIONS OF STAR-LIKE FUNCTIONS VIA CONTINUOUS SEMIGROUPS OF HOLOMOR-PHIC MAPPINGS, Mathematical Proceedings of the Royal Irish Academy, 108 No. 2 (2008) 177–197

(174) A.C. Payatakes - G.N. Constantinides - M.S. Valavanides, Hierarchical theoretical models: An informal introduction, in: G. Dassios (ed.) et al., Mathematical methods in scattering theory and biomedical technology. Proceedings of a workshop dedicated to Professor Gary Roach, Metsovo, Greece, June 30-July 1, 1997. Harlow: Longman. Pitman Res. Notes Math. Ser. 390, 158-169; p. 158 (1998).

(218) D.J. Wales; Energy Landscapes with Applications to Clusters, Biomolecules and Glasses; Cambridge: Cambridge University Press, 2003. 681 p. www.books.google.com; p. 352 (2004).

D.V. Juriev; On the univalency of regular functions; Annali di Mat. Pura ed Appl. 164, No.1, 37–50 (1993).

Revian M. Minyaev - David J. Wales; Gradient line reaction path of HF addition to ethylene; Chem. Phys Letters 218 No.5-6, 403–427 (1994).

K.-H- Indlekofer - N.M. Timofeev; Shifted B-numbers as a set of uniqueness for additive and multiplicative functions; Acta Arithmetica 116 (2005) 295–313.

M.A. Sahai - Sz.-N. Fejér - B. Viskolcz - I. Csizmadia; First principle computational study of the full conformational space of L-theonine diamide, the energetic stability of cis and trans isomeres; J. Phys. Chem. A 110 No.40 (2006) 11527–11536.

W. Kaup (ed.) - C. McCrimmon (ed.) - H. Peterson (ed.); Jordan Algebras; Proceedings of the conference held in Oberwolfach.